# Discovering Global False Negatives On the Fly for Self-supervised Contrastive Learning

### v1

@@ Appendix H Additional Details for Bimodal Experiments
Table 5: Hyperparameters for bimodal training
@@
Appendix I Additional Unimodal Experimental Results
@@
Statistical Significance.
We report in Tables 6 and 7 the p-values with respect to the baseline for the unimodal semi-supervised and transfer learning experiments respectively, and consider a standard significance level of 5%. For the semi-supervised scenario, we can observe GloFND achieves a p-value below 0.018 on all scenarios, thus achieving statistically significant improvements in all scenarios. Moreover, GloFND achieves statistical significance below the 1 % level on both the 100% and 1% scenarios. For the transfer learning case, GloFND achieves statistically significant improvements on 6 out of 8 datasets, while FNC only achieves it on 2 of 8.

### v2

@@ D.2 Additional Details for Bimodal Experiments
Table 4: Hyperparameters for FastCLIP Training
@@
D.3 Additional Experimental Results
Statistical Significance.
We report in Tables 5 and 6 the p-values with respect to the baseline for the unimodal semi-supervised and transfer learning experiments respectively, and consider a standard significance level of 5%. For the semi-supervised scenario, we can observe GloFND achieves a p-value below 0.018 on all scenarios, thus achieving statistically significant improvements in all scenarios. Moreover, GloFND achieves statistical significance below the 1 % level on both the 100% and 1% scenarios. For the transfer learning case, GloFND achieves statistically significant improvements on 6 out of 8 datasets, while FNC only achieves it on 2 of 8.
@@ D.3 Additional Experimental Results
SimCLR Experiment.
We evaluate the performance of GloFND in conjunction with SimCLR. Unlike SogCLR, SimCLR identifies false negatives only within each mini-batch, necessitating the use of a larger batch size. Table 7 presents the linear evaluation results for SimCLR under the experimental setup described in Appendix D.1, using a batch size of 512.
Table 7: Linear evaluation results in unimodal semi-supervised scenario. We train the linear classifiers with different percentages of randomly sampled labeled training data and present their top-1 accuracies (%) on the validation set. We include the overall average and, in parentheses, its improvement WRT SimCLR baseline.
Method 100.0% 10.0% 1.0% 0.1% Average
SimCLR 76.88 73.38 66.40 33.56 62.56
+ FNC 76.90 73.10 64.88 34.20 62.27
+ GloFND 77.14 73.66 66.50 35.58 63.22
CLIP fine-tuning.
To evaluate the effectiveness of GloFND in fine-tuning large pretrained models to mitigate the impact of false negatives, we fine-tune OpenAI’s ResNet-50 CLIP model on CC3M. The experimental setup closely follows Appendix D.2, with the key difference being that we initialize from OpenAI’s pretrained weights and train for 15 epochs, using FNC/GloFND after the first epoch ($\alpha=10^{-4}$). Validation image-text retrieval results are reported in Table 8.
Table 8: We fine-tune OpenAI’s ResNet-50 CLIP model on CC3M and report image-text retrieval results on its validation set.
Method IR@1 IR@5 IR@10 IR Average TR@1 TR@5 TR@10 TR Average
Base 27.58 48.17 57.53 44.43 26.30 48.14 57.69 44.04
GloFND 36.07 58.94 67.22 54.08 35.76 59.19 67.44 54.13
+ FNC 33.69 58.01 67.51 53.07 33.61 57.95 67.53 53.03
+ GloFND 36.52 59.44 68.02 54.66 35.71 59.27 67.97 54.32